OpenStudy (anonymous):

Use spherical coordinates. Evaluate (S = integral sign) SSS(x2 + y2) dV where E lies between the spheres x2 + y2 + z2 = 1 and x2 + y2 + z2 = 25. The answer i came up with was 624.8pi^2

OpenStudy (sidsiddhartha):

ok so what have u done so far? u can show your work

OpenStudy (sidsiddhartha):

start by converting them into spherical coordinates

OpenStudy (sidsiddhartha):

\[x=r.\sin \theta. \cos \phi\\y=r.\sin \theta.\sin \phi\\z=r.\cos \theta\]

OpenStudy (anonymous):

Let me take a picture :)

OpenStudy (sidsiddhartha):

yes the same thing @Concentrationalizing :)

OpenStudy (anonymous):

Just looked odd how ya had it, lol.

OpenStudy (sidsiddhartha):

lol :)

OpenStudy (sidsiddhartha):

try to find the limits first

OpenStudy (anonymous):

\[x = \rho \sin \phi \cos \theta\] \[y = \rho \sin \phi \sin \theta\] \[z = \rho \cos \phi \] had to correct myself :P

OpenStudy (sidsiddhartha):

okey :)

OpenStudy (sidsiddhartha):

\[\int\limits_{0}^{\pi/2}\int\limits_{0}^{\pi/2}\int\limits_{1}^{5}\] seems ok?

OpenStudy (anonymous):

Sorry my picture is REALLY blurry and can't see. I got the limites of integration down. I believe I foudn my mistake give me a second when I tookt he integration for sin^2(phi) i accentally wrote phi as theta..

OpenStudy (anonymous):

May I ask why you limited your phi and theta limits to only pi/2?

OpenStudy (sidsiddhartha):

i think i messed up lol it should be 0 to pi ,and 0 to 2 pi right?

OpenStudy (anonymous):

didnt' see that my limits are 1 <= rho <=5 o <= theta <= 2pi o <= phi <=pi

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Those are the correct limits :3

OpenStudy (sidsiddhartha):

yeah

OpenStudy (sidsiddhartha):

now easy integration

OpenStudy (anonymous):

Right so far?

OpenStudy (anonymous):

Same answer...

OpenStudy (anonymous):

Hmm... \(x^{2} + y^{2} = r^{2}\) \( r= \rho sin \phi\ \implies\ r^{2} = \rho^{2} sin^{2} \phi \implies x^{2} + y^{2} = \rho^{2} sin^{2} \phi \) \[\int\limits_{0}^{2 \pi}\int\limits_{0}^{\pi}\int\limits_{1}^{5} \rho ^{2} \sin^{2} \phi (\rho ^{2} \sin \phi ) d \rho d \phi d \theta\] The \(\rho ^{2} sin ^{2} \phi\) comes from the conversion of the original function \(x^{2} + y^{2} \). The rest comes from the necessary part of an integration using spherical coordinates.

OpenStudy (anonymous):

So I actually messed up on my setup jumped to quick

OpenStudy (sidsiddhartha):

yes did some wrong calculations

OpenStudy (anonymous):

Okay let me do it again don't give me the answer yet :X

OpenStudy (anonymous):

Just combine and make it into \(\rho ^{4} sin^{3} \phi\) and integrate through :)

OpenStudy (sidsiddhartha):

from where P^4 is coming

OpenStudy (anonymous):

Integration in spherical coordinates requires we have \(\rho ^{2} sin \phi\ d \rho\ d \phi\ d \theta\). The conversion of the function inside of the integral gives us: \(x^{2} + y^{2} = p^{2} sin^{2} \phi\) Both of these are being multiplied inside of the integral, giving: \[\int\limits_{0}^{2 \pi}\int\limits_{0}^{\pi} \int\limits_{1}^{5}\rho ^{2} \sin^{2} \phi (p^{2} \sin \phi d \rho d \phi d \theta) = \int\limits_{0}^{2 \pi}\int\limits_{0}^{\pi}\int\limits_{1}^{5} \rho ^{4} \sin^{3} \phi d \rho d \phi d \theta \]

OpenStudy (sidsiddhartha):

yes okayyyy :)

OpenStudy (anonymous):

\[(3124\pi)/5\int\limits_{0}^{2\pi}\sin^3(\phi)dphi\]

OpenStudy (anonymous):

I'm not sure how to handle the \[\sin^3\phi\]

OpenStudy (anonymous):

We can do the half angle identity and get

OpenStudy (anonymous):

\[(sinphi-sinphicos2\phi)/2\]

OpenStudy (anonymous):

Then do a integration by parts?

OpenStudy (anonymous):

\[\sin^{3} \phi = \sin \phi (1 - \cos^{2} \phi) = \sin \phi - \sin \phi \cos^{2} \phi\] Doing this, you can do two integrations now, the integral of \(sin \phi\) alone and a u-substitution in the seciond integral, \(u = cos \phi\)

OpenStudy (sidsiddhartha):

\[\sin^3 \phi=\frac{ 3\sin \phi-\sin3 \phi }{ 4 }\]

OpenStudy (sidsiddhartha):

this will ease ur integration :)

OpenStudy (anonymous):

I don't think I recognize that conversion, can you explain please? It's beautiful btw :')

OpenStudy (anonymous):

Pfft, fancy identities we don't learn, haha. I'll just stick with what I know, lol.

OpenStudy (anonymous):

lol I tried the integration by parts but it just gets me to a loop.

OpenStudy (sidsiddhartha):

its a basic trig indentity \[\sin3A=\sin(2A+A)=\sin2A.cosA+\cos2A.sinA\\=2.sinA.cosA.cosA+(1-2\sin^2A).sinA\\=2sinA(1-\sin^2A)+sinA-2\sin^3A\\=3sinA-4\sin^3A\\sin3A=3sinA-4\sin^3A\\sin^3A=\frac{ 3sinA-\sin3A }{ 4 }\] proved it :)

OpenStudy (anonymous):

But the u sub work s a LOT better..

OpenStudy (sidsiddhartha):

yes its hard to remember them :)

OpenStudy (anonymous):

Yeah, the splitting apart of sin^3 into sin and (1-cos^2) works well. And I remember pretty much any identity that ive learned, but the sin^3 one I havent seen as a basic identity.

OpenStudy (anonymous):

Its always the usual ones, double angle, half angle, power reducing, etc, etc.

OpenStudy (sidsiddhartha):

yes its a important one just makes your calculation easy :)

OpenStudy (anonymous):

So if I were to say \[\sin^4\phi\] = \[(4sinphi-sin4phi)/5\]

OpenStudy (anonymous):

You should end up with a \(sin^{4} \phi\). At least not that I see.

OpenStudy (anonymous):

*shouldnt

OpenStudy (sidsiddhartha):

lol u have to derive it then for sin^n(phi)

OpenStudy (anonymous):

I got the answer is zero.. what did I do wrong -.-

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{\pi}\int\limits_{1}^{5}\rho ^{4} \sin^{3} \phi d \rho d \phi d \theta \] The integration with respect to rho gives \[\frac{ 5^{5} }{ 5 }\sin^{3} \phi - \frac{ 1 }{ 5 }\sin^{3} \phi = \frac{ 3124 }{ 5 }\sin^{3} \phi\] Then we get: \[\int\limits_{0}^{2\pi}\int\limits_{0}^{\pi} \frac{ 3124 }{ 5 }\sin^{3} \phi d \phi d \theta = \int\limits_{0}^{2\pi}\int\limits_{0}^{\pi}\frac{ 3124 }{ 5 }\sin \phi (1- \cos^{2} \phi) d \phi d \theta\] \[\int\limits_{0}^{2\pi}\int\limits_{0}^{\pi}\frac{ 3124 }{ 5 }(\sin \phi - \sin \phi \cos^{2} \phi) d \phi d \theta\] Integrating sin phi is straightforward, integrating the sin cos^2 is just u-sub with u = cos phi \[\frac{ 3124 }{ 5 }(-\cos \phi + \frac{ \cos^{3} \phi }{ 3 })\] THen you have limits from 0 to pi for that integral, then you can do the last integraton with respect to theta. Just to get ya goin :)

OpenStudy (sidsiddhartha):

you've got some concentration @concentrationalizing :P

OpenStudy (anonymous):

:) As long as things don't get too "Complexicated" o_o

OpenStudy (sidsiddhartha):

lol :)

OpenStudy (anonymous):

I'm making so many mistakes stupid rho and phi and rho and asdfasdfhkasgf

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}sinphi = 0\]

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

Because cos(2pi) = 1 and cos(0) = 1 and 1-1 =0

OpenStudy (anonymous):

xDD No worries :) So with where I left off, we have: \[\frac{ 3124 }{ 5 }[(-\cos(\pi)+ \frac{ \cos^{3}(\pi) }{ 3 }) - (-\cos (0) + \frac{ \cos^{3}(0) }{ 3 })]\] \[\frac{ 3124 }{ 5 }[1 - \frac{ 1 }{ 3 }-(-1+ \frac{ 1 }{ 3 })] = \frac{ 3124 }{ 5 }(2 - \frac{ 2 }{ 3 }) = \frac{ 12496 }{ 15 }\] That leavs the final integral to come down to simply \[\int\limits_{0}^{2\pi}\ \frac{ 12496 }{ 15 }d \theta\]

OpenStudy (anonymous):

So 24992pi/15

OpenStudy (anonymous):

Sounds good to me :)

OpenStudy (anonymous):

Can you follow the integration steps, though?

OpenStudy (anonymous):

I figured out my mistake... Looking at my limits of integration I switch phi and theta... so I had 2pi <= phi <= 0

OpenStudy (anonymous):

And yes the steps of integration were perfect thank you, it made it clear to me what my mistake was lol

OpenStudy (anonymous):

I was wondering why you had your limits of integration a bit off, then I realize i was actually the one off.

OpenStudy (anonymous):

Alright, awesome ^_^ Lol, well I make mistakes as well, but I know I dont have any thetas in my integration. And theta is so often 0 to 2pi, so it kind of registers to me to match 0 to pi with phi and 0 to 2pi with theta. That and for the most part, you pretty restrict phi to be between 0 and pi, so not sure, I do my best to not mix stuff up, lol.

OpenStudy (anonymous):

Must not rush these problems again... keeping my limits of integration and what I'm integration with respect is tricky enough lol always feel like I'm being caught in a tangle.

OpenStudy (anonymous):

Thanks! The answer was correct

OpenStudy (anonymous):

Awesome :P Glad I could help :)

OpenStudy (anonymous):

You two are officially my first two fans much appreciated I hope to get some more help you guys later on Exam on Thursday so study study study :)

OpenStudy (anonymous):

Lol, yeah, good luck studying. I got exams on Wednesday o.o So we all study study cram cram now :D